# Towards the Colmez Conjecture

### v1

@@ 1. Introduction
In his 1993 article [Col93], Colmez explicitly constructs a height function for certain periods of a CM abelian variety and conjectures that this height can be given by special values of log derivatives of $L$-functions of Artin characters. One consequence of this conjecture, and how this conjecture is usually presented in the literature, is an explicit formula between the Faltings height of a CM abelian variety and log derivatives of $L$-functions. This conjecture is equivalent to the Chowla–Selberg formula in the case of elliptic curves. We first precisely state the conjecture following [Col93].
@@
Let $\mathbb{Q}^{CM}$ be the maximal CM-extension of $\mathbb{Q}$ and let $c$ denote complex conjugation on $\mathbb{Q}^{CM}$. Let $\mathcal{CM}$ be the set of locally constant functions $f\colon\operatorname{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})\to\mathbb{Q}$ that factor through $\operatorname{Gal}(\mathbb{Q}^{CM}/\mathbb{Q})$ and satisfy the condition that $f(g)+f(cg)$ is independent of $g\in\operatorname{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$. Define $\mathcal{CM}^{0}\subset\mathcal{CM}$ as the subset of class functions, namely those such that $f(ghg^{-1})=f(h)$ for all $g,h\in\operatorname{Gal}(\mathbb{Q}^{CM}/\mathbb{Q})$.
@@
Tensoring up to $\mathbb{C}$, the set of class functions $\mathcal{CM}^{0}\otimes\mathbb{C}$ admits a basis in terms of Artin characters. Let $\mu_{Art}$ be the $\mathbb{C}$-linear functional on $\mathcal{CM}^{0}$ given by $\mu_{Art}(\chi)=\log f_{\chi}$, where $f_{\chi}$ is the conductor of $\chi$, for any Artin character $\chi$. For $s\in\mathbb{C}$, let $Z(\cdot,s)$ be another $\mathbb{C}$-linear functional on $\mathcal{CM}^{0}$ given on Artin characters as $Z(\chi,s)=L^{\prime}(\chi,s)/L(\chi,s)$.
@@
Colmez defines another $\mathbb{C}$-linear functional on $\mathcal{CM}^{0}$ defined by using another basis in terms of CM-types of CM-fields. If $E/\mathbb{Q}$ be a number field, define $H_{E}\coloneqq\operatorname{Hom}(E,\mathbb{C})$. For a CM-field $E$, CM-type $\Phi\subset H_{E}$, and embedding $\tau\in H_{E}$, define the function $a_{\Phi,\tau}$ by the formula
$a_{\Phi,\tau}(g)=\begin{cases}1&\text{if $g\tau\in\Phi$,}\\ 0&\text{otherwise.}\end{cases}$
Let $a_{\Phi,\tau}^{0}\in\mathcal{CM}^{0}$ be its average under conjugation, namely $a_{\Phi,\tau}^{0}=\frac{1}{[K:\mathbb{Q}]}\sum_{\sigma\in H_{K}}a_{\sigma\Phi,\sigma\tau}$ for a number field $K$ containing all conjugates of $E$. Finally set $A_{\Phi}=\sum_{\tau\in\Phi}a_{\Phi,\tau}$ and $A_{\Phi}^{0}=\sum_{\tau\in\Phi}a_{\Phi,\tau}^{0}$.
@@
Let $A$ be an abelian variety with CM of type $(\mathcal{O}_{E},\Phi)$ and let $\tau\in\Phi$. Colmez uses the pairing between $H_{\operatorname{dR}}^{1}(A/\mathbb{C})$ and $H_{1}(A(\mathbb{C}),\mathbb{Z})$ to define a $\tau$-component of the Faltings height, which we calls $ht(\Phi,\tau)$. He then defines $ht(a_{\Phi,\tau}^{0})\coloneqq ht(\Phi,\tau)$ can be extended to a linear height function on $\mathcal{CM}^{0}$. Moreover, it can be related to Faltings height in that if $A$ is a CM abelian variety of type $(\mathcal{O}_{E},\Phi)$, then
@@
Colmez conjectures that the height function given by $ht$ and $Z(\cdot,s)$ are the same.
@@
If $a\in\mathcal{CM}^{0}$ is a class function, let $a^{\vee}\colon\operatorname{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})\to\mathbb{C}$ be the function given by $a^{\vee}(g)=a(g^{-1})$ for all $g\in\operatorname{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$. Then
@@
where $F$ is the maximal totally real subfield of $E$ and $\chi_{E/F}$ is the character associated with the quadratic extension. This averaged version was proven independently by Yuan and Zhang ([YZ18]) and a group of Andreatta, Goren, Howard, and Madapusi ([AGHMP18]). This average result was extended to certain “unitary CM-types of signature $(n-1,1)$” by Yang and Yin ([YY18]). Moreover, using this averaged result, Barquero-Sanchez, Masri, and Thorne were able to prove that $100\%$ of CM abelian varieties satisfy the Conjecture 1.2 ([BSMT23]). This averaged version also has other far-ranging consequences. Tsimerman showed that the averaged Colmez Conjecture implies the André–Oort Conjecture for the moduli space of abelian varieties ([Tsi18]) and later with Pila, Shankar, Esnault, and Groechenig proved that it implied the full André–Oort Conjecture for all Shimura varieties ([PST+24]).
In this article, we prove two main results which serve as steps towards proving Conjecture 1.2 as well as Colmez’s original conjecture that $ht(a)=Z(a^{\vee},s)$ for all $a\in\mathcal{CM}^{0}$. First toward proving Conjecture 1.2, let $\phi\subset\Phi$ be a subset of a CM-type, which we call a partial CM-type. We use the height function $ht$ to propose an explicit definition for the height of a partial CM-type $\phi$. Note that we can simplify $A_{\Phi}$ for a CM-type $\Phi$ as
@@
Let $\phi\subset H_{E}$ be a partial CM-type of $E$ and let $\Sigma\subset H_{F}$ be the restriction of the places in $\phi$ to $F$. Let $A_{\phi}\in\mathcal{CM}$ be given by the formula
$A_{\phi}(g)=\lvert\phi\cap g\phi\rvert+\frac{1}{2}\lvert\Sigma\cap g\Sigma^{c}\rvert.$
@@
Suppose that $E/\mathbb{Q}$ is a Galois CM-extension and if $\sigma,\tau\in H_{E}$ are distinct places so that $\sigma\neq\overline{\tau}$, let $g_{\sigma,\tau}\in\operatorname{Gal}(E/\mathbb{Q})$ be the element such that $g_{0}\sigma=\tau$. Let $c\in\operatorname{Gal}(E/\mathbb{Q})$ denote complex conjugation.
@@
$h(\phi)=$ $-\frac{1}{2}\sum_{\chi}\left(\chi(g_{\sigma,\tau})+\chi(g_{\sigma,\tau}c)+\chi(g_{\sigma,\tau})^{-1}+\chi(g_{\sigma,\tau}c)^{-1}\right)\frac{L^{\prime}(\chi,0)}{L(\chi,0)}$
@@
where $\chi$ runs over all irreducible characters of $\operatorname{Gal}(E/\mathbb{Q})$, the partial CM-type is $\phi=\{\sigma,\tau\}$ with $\Sigma=\{\sigma|_{F},\tau|_{F}\}\subset H_{F}$, and $d_{\phi}$ and $d_{\Sigma}$ are certain root discriminants defined in Section 2.
Then using previous work of the author ([Zha23]), we can restate the formula given in the previous theorem as a period formula for quaternionic Shimura surfaces. We now precisely state this period formula, and refer to [Del71] for more details about Shimura varieties.
Let $E,F,\phi,\Sigma$ be as in the setting of the previous theorem. Let $B/F$ be a quaternion algebra which permits an embedding $E\to B$ and so that $B$ is ramified at infinity exactly at $\Sigma$. Let $G/\mathbb{Q}$ be the algebraic group given by $G\coloneqq\operatorname{Res}_{F/\mathbb{Q}}B^{\times}$. Setting $\mathcal{H}^{\pm}$ to be the upper and lower half-planes, the Shimura datum $(G,(\mathcal{H}^{\pm})^{\Sigma})$ gives a tower of Shimura varieties
@@
The embedding $E\to B$ gives a map $\operatorname{Res}_{E/\mathbb{Q}}\mathbb{G}_{m}\to G$ and the choice of partial CM-type $\phi$ gives us the image of cocharacter of $\operatorname{Res}_{E/\mathbb{Q}}\mathbb{G}_{m}$ in ±$\mathcal{H}^{\pm})^{\Sigma}$ by specifying the upper or lower half-plane, and thus the combination gives a CM-point $P_{U}\subset X_{U}(\overline{\mathbb{Q}})$. An integral model $\mathcal{X}_{U}$ of $X_{U}$ was constructed in [Zha23] and let $\mathcal{P}_{U}$ denote the Zariski closure of $P_{U}$ in $\mathcal{X}_{U}$. Let $h_{\widehat{\mathcal{L}_{U}}}(\mathcal{P}_{U})$ be the Arakelov degree Then a consequence of [Zha23, Thm. 1.1] (and the proof of [Zha23, Thm. 7.5] for the specific log terms) is the equivalence of the following arithmetic period formula with the Colmez Conjecture.
@@
Suppose that $U=\prod_{p}U_{p}$ is a maximal compact subgroup of $G(\mathbb{A}_{f})$. Then Conjecture 1.2 holds for all CM-type of $E$ and all of its CM-subfields if and only if for all choices of $\sigma,\tau$ we have
$\frac{1}{2}h_{\widehat{\mathcal{L}_{U}}}(\mathcal{P}_{U})=$ $-\frac{1}{2}\sum_{\chi}\left(\chi(g_{\sigma,\tau})+\chi(g_{\sigma,\tau}c)+\chi(g_{\sigma,\tau})^{-1}+\chi(g_{\sigma,\tau}c)^{-1}\right)\frac{L^{\prime}(\chi,0)}{L(\chi,0)}$
@@
Then towards Colmez original conjecture, we express his height of CM periods in geometrical terms as the arithmetic degree of a certain line bundle. Namely in [YZ18], for a CM-type $(E,\Phi)$ and $\tau\in\Phi$, they define the $\tau$-component of the height of $h(\Phi)$, which we denote $h(\Phi,\tau)$ as Arakelov degree of a metrized line bundle. We compare this to Colmez’s height function $ht(\Phi,\tau)$ which was defined in terms of period pairings. We prove
@@
$h(\Phi,\tau)=ht(\Phi,\tau)+\frac{1}{2}\log 2\pi+\frac{1}{2[E:\mathbb{Q}]}\log d_{E}-\mu_{Art}(a_{\Phi,\tau}^{0}).$
@@ 1.1. Acknowledgements
I wish to thank Pierre Colmez for encouraging me to propose a definition for a partial CM-type and Shou-Wu Zhang for suggesting that the Colmez Conjecture can be reduced to a $2$-dimensional period formula.

### v2

@@ 1. Introduction
In his 1993 article [COL93], Colmez explicitly constructs a height function for certain periods of a CM abelian variety and conjectures that this height can be expressed in terms of special values of logarithmic derivatives of $L$-functions. One consequence of this conjecture, and how this conjecture is usually presented in the literature, is an explicit formula between the Faltings height of a CM abelian variety and the logarithmic derivatives of $L$-functions. This conjecture is equivalent to the Chowla–Selberg formula in the case of elliptic curves. We first precisely state the conjecture following [COL93].
@@
Let $\mathbb{Q}^{CM}$ be the maximal CM-extension of $\mathbb{Q}$ and let $c$ denote complex conjugation on $\mathbb{Q}^{CM}$. Let $\mathcal{CM}$ be the set of locally constant functions $f\colon\operatorname{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})\to\mathbb{Q}$ that factor through $\operatorname{Gal}(\mathbb{Q}^{CM}/\mathbb{Q})$ and satisfy the condition that $f(x)+f(cx)$ is independent of $x\in\operatorname{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$. Define $\mathcal{CM}^{0}\subset\mathcal{CM}$ as the subset of class functions, namely those functions $f$ such that $f(xyx^{-1})=f(y)$ for all $x,y\in\operatorname{Gal}(\mathbb{Q}^{CM}/\mathbb{Q})$.
@@
Tensoring up to $\mathbb{C}$, the set of class functions $\mathcal{CM}^{0}\otimes\mathbb{C}$ admits a basis in terms of Artin characters. Let $\mu_{Art}$ be the $\mathbb{C}$-linear functional on $\mathcal{CM}^{0}$ given by its value on an Artin character $\chi$ as $\mu_{Art}(\chi)=\log f_{\chi}$, where $f_{\chi}$ is the conductor of $\chi$. For $s\in\mathbb{C}$, let $Z(\cdot,s)$ be another $\mathbb{C}$-linear functional on $\mathcal{CM}^{0}$ given on Artin characters as $Z(\chi,s)=L^{\prime}(\chi,s)/L(\chi,s)$.
@@
Colmez defines another $\mathbb{C}$-linear functional on $\mathcal{CM}^{0}$ defined by using another basis in terms of CM-types of CM-fields. Fix an embedding of $\overline{\mathbb{Q}}\subset\mathbb{C}$. If $E/\mathbb{Q}$ is a number field, define $H_{E}\coloneqq\operatorname{Hom}(E,\overline{\mathbb{Q}})=\operatorname{Hom}(E,\mathbb{C})$. For a CM-field $E$, CM-type $\Phi\subset H_{E}$, and embedding $\tau\in H_{E}$, define the function $a_{\Phi,\tau}\colon\operatorname{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})\to\mathbb{C}$ by the formula
$a_{\Phi,\tau}(x)=\begin{cases}1&\text{if $x\tau\in\Phi$,}\\ 0&\text{otherwise.}\end{cases}$
Let $a_{\Phi,\tau}^{0}\in\mathcal{CM}^{0}$ be its average under conjugation, namely
$a_{\Phi,\tau}^{0}=\frac{1}{[K:\mathbb{Q}]}\sum_{\sigma\in H_{K}}a_{\sigma\Phi,\sigma\tau}$
for a number field $K$ containing all conjugates of $E$. Finally set $A_{\Phi}=\sum_{\tau\in\Phi}a_{\Phi,\tau}$ and $A_{\Phi}^{0}=\sum_{\tau\in\Phi}a_{\Phi,\tau}^{0}$.
@@
Let $A$ be a CM abelian variety of type $(\mathcal{O}_{E},\Phi)$ and let $\tau\in\Phi$. Colmez uses the pairing between $H_{\operatorname{dR}}^{1}(A/\mathbb{C})$ and $H_{1}(A(\mathbb{C}),\mathbb{Z})$ to define a $\tau$-component of the Faltings height, which we denote $ht(\Phi,\tau)$. He then proves that $ht(a_{\Phi,\tau}^{0})\coloneqq ht(\Phi,\tau)$ can be extended to a linear height function, denoted $ht$, on $\mathcal{CM}^{0}$. Moreover, it can be related to the Faltings height of a CM abelian variety. The precise definition of the Faltings height of an abelian variety is given in Definition 3.1.
Theorem 1.1 ([COL93, Conj. II.2.10]).
If $A$ is a CM abelian variety of type $(\mathcal{O}_{E},\Phi)$, then
@@
Since the right hand side depends only on the CM-type $\Phi$, we write $h(\Phi)$ to denote the Faltings height of any (hence all) abelian varieties with CM of type $(\mathcal{O}_{E},\Phi)$.
Colmez conjectures that the height functions given by $ht$ and $Z(\cdot,s)$ are the same.
@@
If $a\in\mathcal{CM}^{0}$ is a class function, let its dual $a^{\vee}\colon\operatorname{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})\to\mathbb{C}$ be the function given by $a^{\vee}(g)=a(g^{-1})$ for all $g\in\operatorname{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$. Then
@@
where $F$ is the maximal totally real subfield of $E$ and $\chi_{E/F}$ is the character associated with the quadratic extension. This averaged version was proven independently by Yuan and Zhang ([YZ18]) and a group of Andreatta, Goren, Howard, and Madapusi ([AGH+18]). This average result was extended to certain “unitary CM-types of signature (n1,1)(n-1,1)” by Yang and Yin ([YY18]). Moreover, using this averaged result, Barquero-Sanchez, Masri, and Thorne were able to prove that $100\%$ of CM abelian varieties satisfy the Conjecture 1.3 ([BMT23]). This averaged version also has other far-ranging consequences. Tsimerman showed that the averaged Colmez Conjecture implies the André–Oort Conjecture for the moduli space of abelian varieties ([TSI18]) and later with Pila, Shankar, Esnault, and Groechenig proved that it implied the full André–Oort Conjecture for all Shimura varieties ([PST+24]).
In this article, we prove two main results which serve as steps towards proving Conjecture 1.3 as well as Colmez’s original conjecture that $ht(a)=Z(a^{\vee},s)$ for all $a\in\mathcal{CM}^{0}$. We define the height of a subset $\phi\subset\Phi$ of a CM-type, which we call a partial CM-type, and express the Faltings height of a CM abelian variety in terms of heights of partial CM-types of order $2$ in Proposition 2.4. A consequence of this result is that Conjecture 1.3 can be reduced to proving explicit identities for heights of partial CM-types of cardinality $2$ (see Theorem 1.5). Combining Theorem 1.5 with [ZHA23] gives Corollary 1.6, which reduces Conjecture 1.3 to an explicit height formula for CM points on a quaternionic Shimura surface. Moreover, we make some progress towards giving a geometric reformulation for the period formula given by Conjecture 1.2. In Theorem 1.7 (also Theorem 4.1), we express the periods given in [COL93] in terms of Arakelov geometry, and by doing so, are able to recover the observation of [YZ18] that the height of nearby CM-types depends only on the CM field, and not on the nearby pair (Corollary 4.2).
We now give a more precise description of our results. For the first result, let $\phi\subset\Phi$ be a partial CM-type. We use the height function $ht$ to propose an explicit definition for the height of a partial CM-type $\phi$. Note that we can simplify $A_{\Phi}$ for a CM-type $\Phi$ as
@@
Let $\phi\subset H_{E}$ be a partial CM-type of $E$ and let $\Sigma\subset H_{F}$ denote the set of places obtained by restricting those in $\phi$ to $F$. Let $A_{\phi}\in\mathcal{CM}$ be given by the formula
$A_{\phi}(x)=\lvert\phi\cap x\phi\rvert+\frac{1}{2}\lvert\Sigma\cap x\Sigma^{c}\rvert$.
@@
Suppose that $E/\mathbb{Q}$ is a Galois CM-extension and if $\sigma,\tau\in H_{E}$ are distinct places so that $\sigma\neq\overline{\tau}$, let $x_{\sigma,\tau}\in\operatorname{Gal}(E/\mathbb{Q})$ be the element such that $x_{\sigma,\tau}\sigma=\tau$. Let $c\in\operatorname{Gal}(E/\mathbb{Q})$ denote complex conjugation.
@@
$h(\phi)=$ $-\frac{1}{2}\sum_{\chi}\left(\chi(x_{\sigma,\tau})+\chi(x_{\sigma,\tau}c)+\chi(x_{\sigma,\tau})^{-1}+\chi(x_{\sigma,\tau}c)^{-1}\right)\frac{L^{\prime}(\chi,0)}{L(\chi,0)}$
@@
where $\chi$ runs over all irreducible characters of $\operatorname{Gal}(E/\mathbb{Q})$, the partial CM-type is $\phi=\{\sigma,\tau\}$ with $\Sigma=\{\sigma|_{F},\tau|_{F}\}\subset H_{F}$, and $d_{\phi}$ and $d_{\Sigma}$ are certain root discriminants defined in Definition 2.7.
Proof.
This immediately follows from Proposition 2.4, Corollary 2.6, and Proposition 2.10. ∎
Using previous work of the author [ZHA23], we can restate the formula given in the previous theorem as a period formula for quaternionic Shimura surfaces. We now precisely state this period formula, and refer to [DEL71] for more details about Shimura varieties.
Let $E,F,\phi,\Sigma$ be as in the setting of the previous theorem. Let $B/F$ be a quaternion algebra which permits an embedding $E\to B$ and assume that at the archimedean places, $B$ is ramified exactly at the infinite places corresponding to $\Sigma$. Let $G/\mathbb{Q}$ be the algebraic group given by $G\coloneqq\operatorname{Res}_{F/\mathbb{Q}}B^{\times}$. Setting $\mathcal{H}^{\pm}$ to be the upper and lower half-planes, the Shimura datum $(G,(\mathcal{H}^{\pm})^{\Sigma})$ gives a tower of Shimura varieties
@@
The embedding $E\to B$ gives a map $\operatorname{Res}_{E/\mathbb{Q}}\mathbb{G}_{m}\to G$ and the choice of partial CM-type $\phi$ gives us the image of cocharacter of $\operatorname{Res}_{E/\mathbb{Q}}\mathbb{G}_{m}$ in $(\mathcal{H}^{\pm})^{\Sigma}$ by specifying the upper or lower half-plane, and thus the combination gives a mapping of Shimura datum and a CM-point $P_{U}\in X_{U}(\overline{\mathbb{Q}})$. An integral model $\mathcal{X}_{U}$ of $X_{U}$ was constructed in [ZHA23] and let $\mathcal{P}_{U}$ denote the Zariski closure of $P_{U}$ in $\mathcal{X}_{U}$. Let $h_{\widehat{\mathcal{L}_{U}}}(\mathcal{P}_{U})$ be the Arakelov degree. Then a consequence of [ZHA23, Thm. 1.1] (and the proof of [ZHA23, Thm. 7.5] for the specific log terms) is the equivalence of the following arithmetic period formula with the Colmez Conjecture.
@@
Suppose that $U=\prod_{p}U_{p}$ is a maximal compact subgroup of $G(\mathbb{A}_{f})$. Then Conjecture 1.3 holds for all CM-types of $E$ and all of its CM-subfields if and only if for all choices of $\sigma,\tau$ we have
$\frac{1}{2}h_{\widehat{\mathcal{L}_{U}}}(\mathcal{P}_{U})=$ $-\frac{1}{2}\sum_{\chi}\left(\chi(x_{\sigma,\tau})+\chi(x_{\sigma,\tau}c)+\frac{1}{\chi(x_{\sigma,\tau})}+\frac{1}{\chi(x_{\sigma,\tau}c)}\right)\frac{L^{\prime}(\chi,0)}{L(\chi,0)}$
@@
Then towards Colmez original conjecture, we express his height of CM periods in geometrical terms as the arithmetic degree of a certain line bundle. Namely in [YZ18], for a CM-type $(E,\Phi)$ and $\tau\in\Phi$, they define the $\tau$-component of the height $h(\Phi)$, which we denote by $h(\Phi,\tau)$, to be the Arakelov degree of a metrized line bundle. We compare this to Colmez’s height function $ht(\Phi,\tau)$ which was defined in terms of period pairings. We prove the following theorem, which appears as Theorem 4.1 later in the article.
@@
Let $E/\mathbb{Q}$ be a CM field with absolute discriminant $d_{E}\in\mathbb{Z}$. Let $\Phi$ be a CM-type of $E$ and $\tau\in\Phi$. Let $h(\Phi,\tau)$ be as defined in [YZ18, Thm. 2.2] and $ht(\Phi,\tau)$ be as defined in [COL93, Lem. II.2.9]. Then
$h(\Phi,\tau)=ht(\Phi,\tau)+\frac{1}{2}\log 2\pi+\frac{1}{2[E:\mathbb{Q}]}\log\lvert d_{E}\rvert-\mu_{Art}(a_{\Phi,\tau}^{0}).$
@@
1.1. Structure of the Article
In Section 2, we prove many results about our height of a partial CM-type. Namely, we express heights of partial CM-types in terms of heights of (full) CM-types in Proposition 2.4 and heights of CM-types in terms of partial CM-types 2.3. Then, we explicitly compute the height of a partial CM-type of size $2$ assuming the Colmez Conjecture in Proposition 2.10. Finally, we give an example showing that these heights of partial CM-types are not comparable in Example 2.11. In Section 3, we recall the definition of the Faltings height of an abelian variety. Moreover, for a CM-type $\Phi$ and $\tau\in\Phi$, we recall the geometric definition of the $\tau$-component of $h(\Phi)$ given in [YZ18, Thm. 2.2], and the definition in terms of periods given in [COL93, Lem. II.2.9]. Then, in Section 4, we give an explicit comparison identity between the geometric and period definitions (Theorem 4.1).
@@ 1.2. Acknowledgements
We wish to thank Pierre Colmez for encouraging me to propose a definition for a partial CM-type and Shou-Wu Zhang for suggesting that the Colmez Conjecture can be reduced to a $2$-dimensional period formula. We would also like to thank Ziqi Guo for asking about how heights of partial CM-types vary for a fixed CM-field $E$. Finally, we would like to thank the anonymous referee for their careful reading and suggestions, improving the presentation of this article.